4. Find each of these values: (a) (177 mod 31 + 270 mod 31) mod 3 (b) (177 mod 31 · 270 mod 31) mod 31
Answer #1

a) 177 mod 31 = 22

270 mod 31 = 22

(177 mod 31 + 270 mod 31) mod 3 = (22 + 22) mod 3 = 44 mod 3 = 2

b) (177 mod 31 - 270 mod 31) mod 3 = (22 - 22) mod 3 = 0 mod 3 = 0